OpenStudy (anonymous):

lim(x+2)/(x^3+8) x-2

OpenStudy (anonymous):

\[\lim \rightarrow 2\]

OpenStudy (anonymous):

This? \[\lim_{x \rightarrow -2} \frac{x+2}{x^3 + 8} \] Gotta use l'Hopitals. \[Let\ g = x+2 \implies g' = 1\] \[Let\ f = x^3+8 \implies f' = 3x^2\] \[\lim_{x \rightarrow -2} \frac{g}{f} = \lim_{x \rightarrow -2} \frac{g'}{f'} = \lim_{x \rightarrow -2}\frac{1}{3x^2} = \frac{1}{12}\]

OpenStudy (anonymous):

yeah thnx a bunch:)